# Better Fair than Sorry: Adversarial Missing Sensitive Value Imputation for Fair GNNs

### v2

@@ 2.3 Proposed Model (BFtS)
Our goal is to learn a fair and accurate GNN $\hat{y}=f_{C}(\mathcal{G},\mathcal{X})$ for node classification with missing sensitive values. The proposed solution is based on adversarial learning and formulates the imputation model as an second adversary of the GNN.
Figure 2 depicts the flow diagram of the proposed model. The model has three primary components: a missing sensitive attribute imputation GNN $f_{I}$, a node classification GNN $f_{C}$, and a sensitive attribute Deep Neural Net(DNN) prediction $f_{A}$. $f_{C}$ takes $\mathcal{X}$ and $\mathcal{G}$ as inputs and predicts the node labels. $f_{A}$ is an adversarial neural network that attempts to estimate the sensitive information from the final layer representations of $f_{C}$ to assess the representations’ bias. More specifically, $f_{C}$ is assumed to be biased if the adversary $f_{A}$ can accurately predict the sensitive attributes information from the representations of $f_{C}$. To train $f_{C}$, we need to know the values of the missing sensitive attributes. $f_{I}$ predicts the missing sensitive attributes. It takes $\mathcal{X}$ and $\mathcal{G}$ as inputs and generates the missing sensitive attributes $\hat{si}$. $f_{I}$ aims to generate sensitive values that minimize the fairness of $f_{C}$. Therefore, it works as a second adversary to the model.
We will first formalize the architecture of the players, then define the their loss functions, and finally describe the training procedure for our approach.
@@ 2.3.1 Player architectures
We describe the architecture of each player in BFtS.
@@
This is a node classification model $\hat{y}_{v}=f_{C}(\mathcal{X}_{v},\mathcal{G})$ implemented using a fair GNN. We assume that $f_{C}$ does not apply sensitive attributes for its training but uses other attributes in $\mathcal{X}$ that might be correlated with sensitive ones. The goal is for $f_{C}$ to be both accurate and fair. To improve fairness, $f_{C}$ tries to minimize the loss of adversary $f_{A}$.
sensitive attribute predictor $f_{A}$:
This is a neural network that applies latent node representations $\mathbf{h}_{v}$ generated by $f_{C}$ to predict sensitive attributes as $\hat{sa}_{v}=f_{A}(\mathbf{h}_{v})$. The idea is that $f_{C}$ is fair if $f_{A}$ performs poorly.
Missing data imputation GNN $f_{I}$:
It predicts missing sensitive attributes as $\hat{si}_{v}=f_{I}(\mathcal{X}_{v},\mathcal{G})$. However, besides being accurate, $f_{I}$ also plays the role of an adversary to $f_{C}$ by predicting values that maximize the accuracy of $f_{A}$.
@@ 2.3.2 Loss Functions
We describe the loss functions applied in our approach. As most real datasets in the fairness literature suffer from data imbalance, we use the Label-Distribution-Aware Margin (LDAM) loss Cao et al. (2019) for classification and imputation.
@@
Let $\hat{y}=f_{C}(X,G)$ be the one hot label predictions and $\hat{y}^{j}_{v}$ be the jth output of node $v$ for the jth class. Then, the LDAM loss of $f_{C}$ is defined as:
$\mathcal{L}_{C}=\frac{1}{|\mathcal{V_{L}}|}\sum_{v=1}^{|\mathcal{V}_{L}|}-log\frac{e^{\hat{y}^{j}_{v}-\Delta_{j}}}{e^{\hat{y}^{j}_{v}-\Delta_{j}+\sum_{k\neq j}e^{\hat{y}^{k}_{v}}}},$ (1)
where $\Delta_{j}=C/n_{j}^{\frac{1}{4}}$. Here $C$ is a hyperparameter and $n_{j}$ is the number of training samples that belong to class $j$.
@@
We also use the LDAM loss to learn the parameters of $f_{I}$:
$\mathcal{L}_{I}=\frac{1}{|\mathcal{V_{S}}|}\sum_{v=1}^{|\mathcal{V}_{S}|}-log\frac{e^{\hat{si}^{j}_{v}-\Delta_{j}}}{e^{\hat{si}^{j}_{v}-\Delta_{j}+\sum_{k\neq j}e^{\hat{y}^{k}_{v}}}}$ (2)
@@
We have the sensitive information of $\mathcal{V}_{s}$. We first replace the $\hat{si}_{v}$ with the $\hat{s}_{v}$ for $v\in\mathcal{V}_{s}$ and generate $\hat{s}$. The parameters of $f_{A}$ are learned using the negative of cross-entropy:
$\mathcal{L}_{A}=$ $\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log f_{A}(\mathbf{h})]$ (3)
$+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{A}(\mathbf{h}))].$
@@ 2.3.3 Learning the parameters of BFtS
$\theta_{C}^{*}=\arg\min_{\theta_{C}}\mathcal{L}_{C}+\alpha\mathcal{L}_{A}$ (4)
where $\alpha$ is a hyperparameter that controls the trade-off between accuracy and fairness of the node classification GNN.
The parameters of the sensitive attribute predictor $\theta_{A}$ only influence the $\mathcal{L_{A}}$ loss. $\theta_{A}$ are learned by simply maximizing the negative of cross-entropy loss:
$\theta_{A}^{*}=\arg\max_{\theta_{A}}\mathcal{L}_{A}$ (5)
Finally, to learn the parameters $\theta_{I}$ of the imputation model toward generating predictions that are accurate and represent the worst-case scenario for the fairness of node representations, we apply the following optimization:
$\theta_{I}^{*}=\arg\min_{\theta_{I}}\mathcal{L}_{I}-\beta\mathcal{L}_{A}$ (6)
Here $\beta$ is a hyperparameter that controls the trade-off between imputation accuracy and the bias generated by the imputed values. The min-max objective between the three players is therefore $\min_{C}\max_{I,A}\mathcal{L}_{A}$.
@@
See the Supp. Material for a pseudocode for BtFS.
VANILLA DEBIAS FAIRGNN RNF BFtS (Ours)
Bail AVPR ($\uparrow$) 0.84±0.00 0.83±0.00 0.83±0.00 0.81±0.00 0.88±0.00
F1 score ($\uparrow$) 0.88±0.00 0.85 ±0.00 0.86±0.00 0.86±0.01 0.91±0.00
%$\%\Delta DP$ ($\downarrow$) 10.10±0.03 8.78±0.02 7.9±0.03 8.50±0.04 7.60±0.03
%$\%\Delta EQOP$ ($\downarrow$) 5.20±0.04 4.50±0.03 4.10±0.04 4.70±0.05 1.70±0.00
Credit AVPR ($\uparrow$) 0.83±0.00 0.82±0.00 0.81±0.00 0.81±0.00 0.81±0.00
F1 score ($\uparrow$) 0.79±0.00 0.73 ±0.00 0.76±0.00 0.75±0.01 0.76±0.00
%$\Delta DP$ ($\downarrow$) 12.10±0.10 6.80±0.20 5.60±0.08 5.30±0.09 5.40±0.02
%$\%\Delta EQOP$ ($\downarrow$) 14.50±0.10 6.20±0.10 5.90±0.10 5.10±0.00 4.50±0.01
German AVPR ($\uparrow$) 0.75±0.00 0.72 ±0.00 0.73±0.00 0.72±0.01 0.72±0.00
F1 score ($\uparrow$) 0.74±0.01 0.73 ±0.00 0.74±0.00 0.72±0.01 0.74±0.00
%$\Delta DP$ ($\downarrow$) 6.30±0.10 5.02±0.20 5.70 ±0.30 4.20±0.00 2.74±0.02
%$\%\Delta EQOP$ ($\downarrow$) 8.40±0.08 6.24±0.20 5.70±0.20 5.90±0.01 1.70±0.04
NBA AVPR ($\uparrow$) 0.75±0.00 0.73±0.00 0.73±0.00 0.73±0.00 0.74±0.00
F1 score ($\uparrow$) 0.73±0.00 0.67 ±0.00 0.70±0.00 0.69±0.00 0.70±0.00
%$\Delta DP$ ($\downarrow$) 8.10±0.02 3.40±0.04 1.20±0.03 1.50±0.05 1.10±0.04
%$\%\Delta EQOP$ ($\downarrow$) 16.50±0.01 5.50±0.06 2.30±0.00 2.10±0.03 1.70±0.03
Table 1: AVPR, F1 score, %$\%\Delta DP$, and %$\%\Delta EQOP$ of different methods. For the Bail, and NBA dataset, we outperform all baselines in terms of accuracy and fairness. For other datasets, we improve fairness while slightly sacrificing accuracy.
@@ 2.4 Theoretical analysis
We will analyze some theoretical properties of BFtS. All the proofs are in the Supp. Material.
Proposition 1.
Given optimal adversaries $f_{A}^{*}$ and $f_{I}^{*}$, the GNN classifier $f_{C}$ solves the following optimization problem:
$\min_{\theta_{C}}\sup_{\hat{s}}JS(p(\mathbf{h}|\hat{s}=1);p(\mathbf{h}|\hat{s}=0)),$ (8)
where $JS$ is the Jensen Shannon divergence.
@@
When the GNN classifier $f_{C}$ reaches global minima, it will achieve demographic parity (Eqn 9) for the worst-case imputation (minimum fairness) $\hat{s}$.
We note that different from Dai and Wang (2021), our guarantee is for the worst-case imputation and not the true sensitive value. That is because our imputations $\hat{s}$ are not independent of the representations $\mathbf{h}$ (see Eqn 7).
Corollary 1.
Let $s^{\prime}=f(G,X)$ be an imputation method independent of models $f_{C}$ and $f_{A}$, then $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))\leq JS(p(\mathbf{h}|\hat{s}=1);p(\mathbf{h}|\hat{s}=0))$.
The value of $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$ is key for the convergence of adversarial learning (see discussion in Supp. Material). For instance, for FairGNN, if $s^{\prime}$ is inaccurate, $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))\approx 0$ and $\mathcal{L}_{A}$ in Eqn 4 will be a constant. The corollary shows that our method is more robust to this issue—as also illustrated in our experiments.
3 Experimental Results
Figure 3: Performance of the methods using the Simulation dataset for different values of assortativity coefficients. In the x-axis, we plot the F1 score and in the y axis of the top row we plot $1-DP$ and in the y-axis of the top row we plot $1-EQOP$. The top right corner of the plot, therefore, represents a high F1 with low bias. When the assortativity is low, other methods fail to learn the node labels. With higher assortativity, though other methods learn the class labels, BFtS is less biased for similar accuracy. Note that the X-axes have different ranges.
We compare our approach for learning fair GNNs for node classification (BtFS) against the state-of-the-art in terms of accuracy and fairness using real and synthetic datasets.
@@ 3 Experimental Results
Figure 5: 2-D kernel PCA node representations for the Bail dataset generated by the FairGNN and BFtS (our approach). We use different markers for nodes depending on their predicted class, sensitive attribute (race), and whether the sensitive attribute is missing. The results show that representations for nodes with missing values are more biased for FairGNN—they are concentrated in a negative class region. On the other hand, nodes with missing values are better spread over the space for BFtS representations. In BFtS, there are more ”Race = Black” nodes than in FairGNN that have missing sensitive values in training but are predicted to be positive.
3.1 Datasets
The datasets are described in the Supp. Materials.
3.2 Baselines
Vanilla: The vanilla model is the basic GNN model trained with cross-entropy loss without any fairness intervention.
Debias Zhang et al. (2018): Debias is an adversarial learning method that trains a classifier and an adversary simultaneously. The adversary is trained using the softmax output of the last layer of the classifier. It does not handle missing sensitive data. Hence, we only use $\mathcal{V}_{S}\in\mathcal{V}$ to calculate the adversarial loss.
FairGNN Dai and Wang (2021): FairGNN first uses a GNN to estimate the missing sensitive attributes. It then trains another GNN-based classifier and a DNN adversary. The adversary helps to eliminate bias from the representations learned by the GNN classifier.
RNF Du et al. (2021): RNF imputes the missing sensitive attributes by first training the model with generalized cross entropy Zhang and Sabuncu (2018). It eliminates bias from the classification head. RNF uses samples with the same ground-truth label but distinct sensitive attributes and trains the classification head using their neutralized representations.

### v3

@@ 3.1 Proposed Model (BFtS)
Our goal is to learn a fair and accurate GNN $\hat{y}=f_{C}(\mathcal{G},\mathcal{X})$ for node classification with missing sensitive values. The proposed solution is model-agnostic and based on adversarial learning. It formulates the imputation model as a second adversary of the GNN.
Figure 3 depicts the flow diagram of the proposed model. The model has three primary components: a missing sensitive attribute imputation GNN $f_{I}$, a node classification GNN $f_{C}$, and a sensitive attribute Deep Neural Net (DNN) prediction $f_{A}$. $f_{C}$ takes $\mathcal{X}$ and $\mathcal{G}$ as inputs and predicts the node labels. $f_{A}$ is an adversarial neural network that attempts to estimate the sensitive information from the final layer representations of $f_{C}$ to assess the bias. More specifically, it $f_{C}$ is assumed to be biased if the adversary $f_{A}$ can accurately predict the sensitive attribute information from the representations of $f_{C}$. To train $f_{C}$, we need to know the values of the missing sensitive attributes. $f_{I}$ predicts the missing sensitive attributes. It takes $\mathcal{X}$ and $\mathcal{G}$ as inputs and generates the missing sensitive attributes $\hat{si}$. $f_{I}$ aims to generate sensitive values that minimize the fairness of $f_{C}$. Therefore, it works as a second adversary to the model.
We will formalize the architecture of the players, define their loss functions, and describe the training procedure for our approach in the next section.
@@ 3.1.1 Player architectures
This section describes the architecture of each player in BFtS.
@@
Node classification model $\hat{y}_{v}=f_{C}(x_{v},\mathcal{G})$ implemented using a GNN. We assume that $f_{C}$ does not apply sensitive attributes but uses other attributes in $\mathcal{X}$ that might be correlated with sensitive ones. The goal of $f_{C}$ is to achieve both accuracy and fairness. To improve fairness, $f_{C}$ tries to minimize the loss of adversary $f_{A}$.
Sensitive attribute predictor (Adversary 1) $f_{A}$:
Neural network that applies representations $\mathbf{h}_{v}$ generated by $f_{C}$ to predict sensitive attributes as $\hat{sa}_{v}=f_{A}(\mathbf{h}_{v})$. $f_{C}$ is fair if $f_{A}$ performs poorly.
Missing data imputation GNN (Adversary 2) $f_{I}$:
Predicts missing sensitive attributes as $\hat{si}_{v}=f_{I}(x_{v},\mathcal{G})$. However, besides being accurate, $f_{I}$ plays the role of an adversary to $f_{C}$ by predicting values that maximize the accuracy of $f_{A}$.
@@ 3.1.2 Loss Functions
Here, we describe the loss functions applied in our approach.
@@
Let $\hat{y}=f_{C}(X,G)$ be the predictions of $f_{C}$. We apply the cross-entropy loss to $f_{C}$’s predictions as follows:
$\mathcal{L}_{C}=-\frac{1}{|\mathcal{V_{L}}|}\sum_{v\in\mathcal{V_{L}}}\hat{y}_{v}log(\hat{y}_{v})+(1-\hat{y}_{v})log(1-\hat{y}_{v}).$ (1)
@@
Because sensitive attributes tend to be imbalanced in real datasets, $f_{I}$ applies the Label-Distribution-Aware Margin (LDAM) loss Cao et al. (2019) for classification. Let $\hat{si}=f_{I}(X,G)$ be the one-hot encoded predictions and $\hat{si}^{s}_{v}$ be the output of node $v$ for $s_{v}$. Then, the LDAM loss of $f_{I}$ is defined as:
$\mathcal{L}_{I}=\frac{1}{|\mathcal{V_{S}}|}\sum_{v=1}^{|\mathcal{V}_{S}|}-log\frac{e^{\hat{si}^{s}_{v}-\Delta^{s}}}{e^{\hat{si}^{s}_{v}-\Delta^{s}}+\sum_{k\neq s}e^{\hat{si}^{k}_{v}}}.$ (2)
where $\Delta^{j}=C/n_{j}^{\frac{1}{4}}$ and $j\in\{0,1\}$.
Here $C$ is a constant independent of the sensitive attribute $s$, and $n_{j}$ is the number of samples that belong to $s=j$. The LDAM loss is a weighted version of the negative log-likelihood loss. Intuitively, since $\Delta^{j}$ is larger for smaller values of $n_{j}$, it ensures a higher margin for the smaller classes.
Furthermore, BFtS can generate fair outputs when very little or no sensitive information is provided. When no or very little sensitive information is available, ($|\mathcal{V}_{S}|\to 0$), $f_{I}$ can impute the sensitive values from the ground truth label $y$ with LDAM loss. We can replace $\mathcal{V}_{S}$ with $\mathcal{V}_{L}$ and $s$ with $y$ for each node $v\in\mathcal{V}_{L}$ in Eqtn 2. The primary argument is that the worst-case fairness model generally assigns a more desired outcome to the non-protected group and a less desired outcome to the protected group. Therefore, the nodes predicted as the minority class will fall into the protected group, and vice versa. Compared to the model $f_{I}$ trained with LDAM loss with limited sensitive information of $\mathcal{V}_{S}$; without any sensitive information, $f_{I}$ trained with LDAM loss learns more bias-amplified samples with labels $y$ of $V_{L}$. Depending on the amount of sensitive information available, the user can choose any of these two methods to learn the imputation.
@@
Given the sensitive information for $\mathcal{V}_{s}$, we first replace $\hat{si}_{v}$ by ${s}_{v}$ for $v\in\mathcal{V}_{s}$ and thereby generate $\hat{s}$. The parameters of $f_{A}$ are learned using:
$\mathcal{L}_{A}=$ $\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log f_{A}(\mathbf{h})]$
$+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{A}(\mathbf{h}))]$ (3)
@@ 3.1.3 Learning the parameters of BFtS
$\theta_{C}^{*}=\arg\min_{\theta_{C}}\mathcal{L}_{C}+\alpha\mathcal{L}_{A}.$ (4)
where $\alpha$ is a hyperparameter that controls the trade-off between accuracy and fairness.
The parameters of the sensitive attribute predictor $\theta_{A}$ only influence the $\mathcal{L_{A}}$ loss. $\theta_{A}$ are learned by maximizing $\mathcal{L}_{A}$:
$\theta_{A}^{*}=\arg\max_{\theta_{A}}\mathcal{L}_{A}.$ (5)
To learn the parameters $\theta_{I}$ of the imputation model to generate predictions that are accurate and represent the worst-case scenario for fairness, we apply the following:
$\theta_{I}^{*}=\arg\min_{\theta_{I}}\mathcal{L}_{I}-\beta\mathcal{L}_{A}.$ (6)
Here $\beta$ is a hyperparameter that controls the trade-off between imputation accuracy and worst-case imputation. The min-max objective between the three players is therefore:
@@
3.2 Training Algorithm
Algorithm 1 is a high-level description of the key steps applied for training BFtS. It receives the graph $\mathcal{G}$, labels $y$, labeled nodes $\mathcal{V}_{L}$, nodes with observed protected attributes $\mathcal{V}_{S}$ and hyperparameters, $\alpha$ and $\beta$ as inputs and outputs the GNN classifier $f_{C}$, sensitive attribute predictor, $f_{A}$ and missing data imputation GNN $f_{I}$. We first estimate $\hat{si}$ and then fix $\theta_{C}$ and $\theta_{A}$ to update $\theta_{I}$. Then we update $\theta_{A}$ with Eq. 5. After that, we fix $\theta_{A}$ and $\theta_{I}$ and update $\theta_{C}$ with Eq. 4 and repeat these steps until convergence. The time complexity analysis of the algorithm is discussed in the Supp. material.
Algorithm 1 Training BFtS
Input: $\mathcal{G}$; $y$; $\mathcal{V}_{L}$; $\mathcal{V}_{S}$; $\alpha$, $\beta$
Output: $f_{C}$, $f_{A}$, $f_{I}$
1:repeat
2:     Get the estimated sensitive attributes $\hat{si}$ with $f_{I}$
3:     Fix $\theta_{C}$ and $\theta_{A}$ and update $\theta_{I}$ using Eq. 6
4:     Update $\theta_{A}$ using Eq. 5
5:     Fix $\theta_{I}$ and $\theta_{A}$ and update $\theta_{C}$ using Eq. 4
6:until converge
@@ 3.3 Theoretical analysis
We will analyze some theoretical properties of BFtS. All the proofs are provided in the Supp. material.
@@
The optimal $f^{*}_{I}$ generates the worst-case (minimum fairness) imputation so that $\Delta DP_{f_{C},f^{*}_{I}}\geq\Delta DP_{f_{C},f_{I}}$ where, $\Delta DP_{f_{C},f_{I}}=|p(\hat{y}=1|\hat{s}=1)-p(\hat{y}=1|\hat{s}=0)|$, $\hat{y}=f_{C}(\mathcal{G},\mathcal{X})$ and, $\hat{s}=f_{I}(\mathcal{G},\mathcal{X}).$
$f_{I}$ generates imputations such that $\Delta DP$ in Eqn 8 is maximized for a given classifier $f_{C}$
Theorem 2.
$f_{C}$ is a minimax estimator, its maximal $\Delta DP$ is minimal among all estimators of $s$.
The optimal GNN classifier $f^{*}_{C}$, will achieve demographic parity ($\Delta DP=0$ in Eqn 8) for the worst-case imputation (minimum fairness) $\hat{s}$ generated by $f_{I}$.
Corollary 2.1.
Let $s^{\prime}=f(G,X)$ be an imputation method independent of models $f_{C}$ and $f_{A}$, then the Jensen Shannon divergence $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))\leq JS(p(\mathbf{h}|\hat{s}=1);p(\mathbf{h}|\hat{s}=0))$.
The value of $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$ is related to the convergence of adversarial learning. For FairGNN, if $s^{\prime}$ is inaccurate, $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))\approx 0$ and $\mathcal{L}_{A}$ in Eq. 4 will be constant. The interplay between the three players makes BFtS more robust to convergence issues because the objective minimizes the upper bound on the JS divergence. This reduces the probability that the divergence vanishes during training (see details in Supp. material).
4 Experiments and Results
Figure 4: Performance of the methods using the Simulation dataset for different values of assortativity coefficients. In the x-axis, we plot the F1 score, and in the y-axis of the top row, we plot $1-\Delta DP$, and in the y-axis of the bottom row, we plot $1-\Delta EQOP$. The top right corner of the plot, therefore, represents a high F1 with low bias. When the assortativity is low, other methods fail to learn the node labels. With higher assortativity, though other methods learn the class labels, BFtS is less biased and has similar accuracy. Note that the X-axes have different ranges.
We compare our approach for learning fair GNNs for node classification (BtFS) against the state-of-the-art in terms of accuracy and fairness using real and synthetic data. Our code can be found in an anonymous repository [^1] https://anonymous.4open.science/r/BFtS-6ADA. The details of the datasets, baselines, and hyperparameters are in the Supp. material. Results with graph attention network and ablation study are also in the Supp. material.
@@ 4 Experiments and Results
RNF BFtS
AVPR($\uparrow$) F1 Score($\uparrow$) %$\%\Delta DP$($\downarrow$) %$\%\Delta EQOP(\downarrow$) AVPR($\uparrow$) F1 Score($\uparrow$) %$\%\Delta DP$ ($\downarrow$) %$\%\Delta EQOP(\downarrow)$
Bail 0.81±0.00 0.85±0.00 11.65±0.02 8.51±0.02 0.83±0.00 0.86±0.00 8.01±0.03 4.12±0.01
Credit 0.80±0.00 0.76±0.00 8.01±0.08 7.25±0.08 0.82±0.00 0.77±0.00 5.97±0.10 4.96±0.06
German 0.73±0.00 0.74±0.00 9.42±0.04 8.92±0.12 0.74±0.00 0.75±0.00 6.42±0.11 7.68±0.17
NBA 0.70±0.00 0.70±0.00 6.47±0.03 5.89±0.05 0.73±0.00 0.70±0.00 5.19±0.01 3.59±0.05
Pokec-z 0.73±0.00 0.71±0.00 6.18±0.05 6.29±0.07 0.73±0.00 0.73±0.00 5.10±0.02 3.20±0.06
Pokec-n 0.72±0.00 0.72±0.00 7.58±0.01 7.09±0.04 0.73±0.00 0.70±0.00 4.29±0.02 3.01±0.03
Table 1: AVPR, F1 score, %$\%\Delta DP$, and %$\%\Delta EQOP$ without any protected information of BFtS and RNF. Amongst the baselines, only RNF works without any protected information. BFtS outperforms RNF in terms of fairness and accuracy for all baselines.
4.1 Selection of Missing Value
We generate missing values synthetically in our experiments. To simulate an adversarial missingness process, we select low-degree nodes to have missing values. This choice is based on previous work showing that low-degree nodes have lower prediction accuracy Tang et al. (2020); Liu et al. (2023); Ju et al. (2024). To generate $k$ nodes with missing sensitive data, we choose the $k$ nodes with the lowest degree. The size of $\mathcal{V}_{S}$ was set to $30\%$ of $|\mathcal{V}|$ unless stated otherwise.